Bakery is looking into a new type of bag tie. According to a competitive analysis of Juan & Juana bread’s compared to those of the competition, the new bag ties would dramatically increase the shelf life of the bread. Juan & Juana bakery currently sells all the bread the company can make, with a gross margin of Php 15,000 a week. The new tie would cost Php 2.00 more than the old one, but the longer shelf life would create incremental value to the customer. Assume that the new tie would require a machine that would add Php 4,000 per week to the company's fixed costs. How much more could Juan & Juana bakery charge per bag in order to make the same gross margin as before?
1
Expert's answer
2019-09-03T10:15:44-0400

P*Q - TC = 15,000.

If gross margin is the same as before, then:

P*Q - (AVC*Q + FC) = (P + x)*Q - ((AVC + 2)*Q + FC + 4000),

P*Q - AVC*Q - FC = P*Q + x*Q - AVC*Q - 2Q - FC - 4000,

x*Q - 2Q = 4000,

x = 4000/Q + 2, where Q is the quantity of bread sold.